# Decision making in stochastic extensive form I: Stochastic decision forests

### v1

@@ A.2 Section 2
Let $\bar{\bm{\mathsf{X}}}={\bm{\mathsf{X}}}\setminus\{\emptyset\}$, that is the set of all random moves ${\bm{\mathsf{x}}}$ with $D_{\bm{\mathsf{x}}}\neq\emptyset$. Then it is evident that $\bar{\bm{\mathsf{X}}}$ satisfies properties 3a, 3b, 3c and 3d of definition 2.9, and is refined by ${\bm{\mathsf{X}}}$. Hence, by the maximality property 3e in definition 2.9, $\bar{\bm{\mathsf{X}}}={\bm{\mathsf{X}}}$. ∎
@@
(Ad “order embedding”):  Let $({\bm{\mathsf{y}}}_{1},\omega_{1}),({\bm{\mathsf{y}}}_{2},\omega_{2})\in{\bm{\mathsf{T}}}\bullet\Omega$. If both ${\bm{\mathsf{y}}}_{1},{\bm{\mathsf{y}}}_{2}$ are elements of ${\bm{\mathsf{X}}}$, the implication $\Rightarrow$” is a consequence of the definition of $\geq_{\bm{\mathsf{X}}}$. The implication $\Leftarrow$” can be shown as follows: If ${\bm{\mathsf{y}}}_{1}(\omega_{1})\supseteq{\bm{\mathsf{y}}}_{2}(\omega_{2})$, then both moves belong to the same connected component of the rooted forest $(F,\supseteq)$ (whose existence axiom 1 ensures). By axioms 2 and 3a, $\omega_{1}=\omega_{2}$. Then, axiom 3c implies that ${\bm{\mathsf{y}}}_{1}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{y}}}_{2}$, whence ${\bm{\mathsf{y}}}_{1}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{2}$.
@@
If $k=2$, then ${\bm{\mathsf{y}}}_{1}(\omega_{1})\supseteq{\bm{\mathsf{y}}}_{2}(\omega_{2})$ is equivalent to the conjunction of these three statements: $\omega_{1}=\omega_{2}$, $\omega_{2}\in D_{{\bm{\mathsf{y}}}_{1}}$, and $w_{k}\in{\bm{\mathsf{y}}}_{1}(\omega_{1})$. By definition of $\geq_{\bm{\mathsf{T}}}$, this is equivalent to ${\bm{\mathsf{y}}}_{1}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{2}$ and $\omega_{1}=\omega_{2}$.
(Ad “bijection”):  Let $({\bm{\mathsf{y}}}_{1},\omega_{1}),({\bm{\mathsf{y}}}_{2},\omega_{2})\in{\bm{\mathsf{T}}}\bullet\Omega$ such that ${\bm{\mathsf{y}}}_{1}(\omega_{1})={\bm{\mathsf{y}}}_{2}(\omega_{2})$. As the evaluation map is an order embedding, which we have proven just before, we obtain ${\bm{\mathsf{y}}}_{1}={\bm{\mathsf{y}}}_{2}$ and $\omega_{1}=\omega_{2}$. Regarding surjectivity, let $y\in F$. If $y$ is terminal, then there is $w\in W$ with $y=\{w\}$ because $F$ is a decision forest over $W$ [^16] Although this is discussed in [3], we give an argument here for the reader’s convenience. Let $f$ be the map from definition 1.3. If $y$ is terminal, then $(\mathcal{P}f)(y)=W(y)$ contains exactly one element, namely $\uparrow y$. As $f$ is injective, $y$ must be a singleton. Hence, $y={\bm{\mathsf{y}}}(\omega)$ for $\omega=\pi(y)$ and the random terminal node ${\bm{\mathsf{y}}}=(\omega,w)$. If $y$ is a move, then, by the definition of ${\bm{\mathsf{X}}}$, property 3b, there is ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ and $\omega\in D_{\bm{\mathsf{x}}}$ such that ${\bm{\mathsf{x}}}(\omega)=y$. ∎
@@
(Ad “forest”):  Let ${\bm{\mathsf{y}}}_{0}\in{\bm{\mathsf{T}}}$ and ${\bm{\mathsf{y}}}_{1},{\bm{\mathsf{y}}}_{2}\in\uparrow{\bm{\mathsf{y}}}_{0}$. As $D_{{\bm{\mathsf{y}}}_{0}}\neq\emptyset$ by lemma 2.10 and the definition of random terminal nodes, there is $\omega\in D_{{\bm{\mathsf{y}}}_{0}}$, and we have $\omega\in D_{{\bm{\mathsf{y}}}_{k}}$ and ${\bm{\mathsf{y}}}_{k}(\omega)\supseteq{\bm{\mathsf{y}}}_{0}(\omega)$ for both $k=1,2$. As $(F,\supseteq)$ is a forest, $\uparrow{\bm{\mathsf{y}}}_{0}(\omega)$ is a chain, thus ${\bm{\mathsf{y}}}_{1}(\omega)\supseteq{\bm{\mathsf{y}}}_{2}(\omega)$ or ${\bm{\mathsf{y}}}_{2}(\omega)\supseteq{\bm{\mathsf{y}}}_{1}(\omega)$. Hence, by proposition 2.11, we get ${\bm{\mathsf{y}}}_{1}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{2}$ or ${\bm{\mathsf{y}}}_{2}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{1}$. We conclude that $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ is a forest.
@@
(Ad “decision forest”):  Let ${\bm{\mathsf{y}}}_{1},{\bm{\mathsf{y}}}_{2}\in{\bm{\mathsf{T}}}$ be such that ${\bm{\mathsf{y}}}_{1}\neq{\bm{\mathsf{y}}}_{2}$. If they are not comparable by $\geq_{\bm{\mathsf{T}}}$, that is neither ${\bm{\mathsf{y}}}_{1}>_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{2}$ nor ${\bm{\mathsf{y}}}_{2}>_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{1}$, then any maximal chain ${\bm{\mathsf{w}}}$ in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ with ${\bm{\mathsf{y}}}_{1}\in{\bm{\mathsf{w}}}$ satisfies ${\bm{\mathsf{y}}}_{2}\notin{\bm{\mathsf{w}}}$. Using the axiom of choice in the form of the Hausdorff maximal principle, there exists indeed such a ${\bm{\mathsf{w}}}$.
@@
For symmetry reasons, it remains to consider the case in that we have ${\bm{\mathsf{y}}}_{1}>_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{2}$. Then, by proposition 2.11, we have for all $\omega^{\prime}\in D_{{\bm{\mathsf{y}}}_{2}}$, ${\bm{\mathsf{y}}}_{1}(\omega^{\prime})\supsetneq{\bm{\mathsf{y}}}_{2}(\omega^{\prime})$. As $D_{{\bm{\mathsf{y}}}_{2}}\neq\emptyset$, we can select $\omega\in D_{{\bm{\mathsf{y}}}_{2}}$. We can also select $w\in{\bm{\mathsf{y}}}_{1}(\omega)\setminus{\bm{\mathsf{y}}}_{2}(\omega)$. In particular, ${\bm{\mathsf{y}}}_{2}(\omega)\notin\uparrow\{w\}$. As the latter set is a maximal chain in $(F,\supseteq)$, there is $y_{3}\in\uparrow\{w\}$ such $y_{3}$ and ${\bm{\mathsf{y}}}_{2}(\omega)$ cannot be compared by $\supseteq$. As ${\bm{\mathsf{y}}}_{1}(\omega)$ is also an element of the chain $\uparrow\{w\}$, this implies ${\bm{\mathsf{y}}}_{1}(\omega)\supsetneq y_{3}$. By proposition 2.11, there is ${\bm{\mathsf{y}}}_{3}\in{\bm{\mathsf{T}}}$ such that $\omega\in D_{{\bm{\mathsf{y}}}_{3}}$ and $y_{3}={\bm{\mathsf{y}}}_{3}(\omega)$. This lemma also implies that ${\bm{\mathsf{y}}}_{2}$ and ${\bm{\mathsf{y}}}_{3}$ are not comparable via $\geq_{\bm{\mathsf{T}}}$. Hence, as we have proven just above, there is a maximal chain ${\bm{\mathsf{w}}}$ in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ with ${\bm{\mathsf{y}}}_{2}\notin{\bm{\mathsf{w}}}$ and ${\bm{\mathsf{y}}}_{3}\in{\bm{\mathsf{w}}}$. By proposition 2.11, we obtain ${\bm{\mathsf{y}}}_{1}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{3}$, and using the fact that $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ is a forest, as already proven, we infer that ${\bm{\mathsf{y}}}_{1}\in{\bm{\mathsf{w}}}$. Hence there is a maximal chain in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ separating ${\bm{\mathsf{y}}}_{1}$ and ${\bm{\mathsf{y}}}_{2}$.
(Ad “rooted”):  Suppose that $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ were not rooted. As $(F,\supseteq)$ is rooted and all roots are moves, by axioms 3b and 3d there is at least one random move whose image consists of roots. If $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ were not rooted, there would have to be at least two of them. Hence, there would be ${\bm{\mathsf{x}}}_{0},{\bm{\mathsf{x}}}_{1}\in{\bm{\mathsf{X}}}$ whose images are roots of $F$ and such that $D_{{\bm{\mathsf{x}}}_{0}}\cap D_{{\bm{\mathsf{x}}}_{1}}=\emptyset$, by axiom 3c. Then let $\bar{\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{0}\cup{\bm{\mathsf{x}}}_{1}$, [^17] That is, $\bar{\bm{\mathsf{x}}}\colon D_{{\bm{\mathsf{x}}}_{0}}\cup D_{{\bm{\mathsf{x}}}_{1}}\to X$ such that $\bar{\bm{\mathsf{x}}}(\omega)={\bm{\mathsf{x}}}_{k}(\omega)$ for $\omega\in D_{{\bm{\mathsf{x}}}_{k}}$, for both $k=0,1$. and $\bar{\bm{\mathsf{X}}}=({\bm{\mathsf{X}}}\setminus\{{\bm{\mathsf{x}}}_{0},{\bm{\mathsf{x}}}_{1}\})\cup\{\bar{\bm{\mathsf{x}}}\}$.
Clearly, $D_{\bar{\bm{\mathsf{x}}}}=D_{{\bm{\mathsf{x}}}_{0}}\cup D_{{\bm{\mathsf{x}}}_{1}}$ would be an element of $\mathscr{A}$ and $\bar{\bm{\mathsf{x}}}$ would be a section of $\pi$ on $D_{\bar{\bm{\mathsf{x}}}}$. It is evident that $\bar{\bm{\mathsf{X}}}$ would induce a covering of $X$ and identify roots in the sense of axiom 3d. To show that $\bar{\bm{\mathsf{X}}}$ would satisfy axiom 3c, it suffices to show that any ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}\setminus\{{\bm{\mathsf{x}}}_{0},{\bm{\mathsf{x}}}_{1}\}$ admitting $\omega\in D_{\bm{\mathsf{x}}}\cap D_{\bar{\bm{\mathsf{x}}}}$ with $\bar{\bm{\mathsf{x}}}(\omega)\supseteq{\bm{\mathsf{x}}}(\omega)$ would have to satisfy $\bar{\bm{\mathsf{x}}}\geq_{\bar{\bm{\mathsf{X}}}}{\bm{\mathsf{x}}}$. But this follows easily because the hypothesis on ${\bm{\mathsf{x}}}$ and $\omega$ implies that there is $k=0,1$ such that $\omega\in D_{{\bm{\mathsf{x}}}}\cap D_{{\bm{\mathsf{x}}}_{k}}$ and ${\bm{\mathsf{x}}}_{k}(\omega)\supseteq{\bm{\mathsf{x}}}(\omega)$. Hence, ${\bm{\mathsf{x}}}_{k}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}$ which directly yields $\bar{\bm{\mathsf{x}}}\geq_{\bar{\bm{\mathsf{X}}}}{\bm{\mathsf{x}}}$. By construction, ${\bm{\mathsf{X}}}$ would refine $\bar{\bm{\mathsf{X}}}$. Thus, axiom 3e would imply the false statement $\bar{\bm{\mathsf{X}}}={\bm{\mathsf{X}}}$.
We conclude that $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ must be rooted.
@@
(Ad ${\bm{\mathsf{X}}}$ = set of moves):  Let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$. Then there is $\omega\in D_{\bm{\mathsf{x}}}$ and, by axiom 3a, ${\bm{\mathsf{x}}}(\omega)\in X$. Hence, there is $y\in F$ with ${\bm{\mathsf{x}}}(\omega)\supsetneq y$. By proposition 2.11, there is unique ${\bm{\mathsf{y}}}\in{\bm{\mathsf{T}}}$ with $\omega\in D_{\bm{\mathsf{y}}}$ and $y={\bm{\mathsf{y}}}(\omega)$, and whence, by the same proposition, ${\bm{\mathsf{x}}}>_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}$. Hence, ${\bm{\mathsf{x}}}$ is a move in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$.
@@
On the other hand, the elements of ${\bm{\mathsf{T}}}\setminus{\bm{\mathsf{X}}}$, i.e., the random terminal nodes, are terminal nodes in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ directly by definition of $\geq_{\bm{\mathsf{T}}}$. Hence, ${\bm{\mathsf{X}}}$ is the set of moves in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$. ∎
@@
$(\Omega,\mathscr{A})$ is a measurable space. $F$ is clearly a $W$-poset and the map assigning to any $w\in W$ the chain $\uparrow\{w\}$ is easily seen to be a bijection between $W$ and the set of maximal chains in $(F,\supseteq)$. Moreover, any chain in $(F,\supseteq)$ contains a maximal element, hence, by [3, theorem 3], expressed in the language of the present paper, $F$ a decision forest over $W$ (axiom 1).
@@
For all $k=0,1,2$, ${\bm{\mathsf{x}}}_{k}$ is defined on $\Omega$ which is an event. Clearly, ${\bm{\mathsf{x}}}_{k}$ is a section of $\pi$. The union of the images of ${\bm{\mathsf{X}}}$’s elements is the set of nodes with at least two elements. As $\{\{w\}\mid w\in W\}\subseteq F$, this union equals $X$. Thus axioms 3a and 3b are satisfied.

### v2

@@ A.2 Section 2
(Ad Part 1):  Suppose that $(F,\pi,{\bm{\mathsf{X}}})$ is order consistent. Let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$. Suppose there is $\omega\in D_{\bm{\mathsf{x}}}$ such that ${\bm{\mathsf{x}}}(\omega)=W_{\omega}$. Let $\omega^{\prime}\in D_{\bm{\mathsf{x}}}$. It suffices to show that ${\bm{\mathsf{x}}}(\omega^{\prime})=W_{\omega^{\prime}}$.
By the covering property in Axiom 33(b), there is ${\bm{\mathsf{x}}}^{\prime}\in{\bm{\mathsf{X}}}$ such that ${\bm{\mathsf{x}}}^{\prime}(\omega^{\prime})=W_{\omega^{\prime}}$. Then ${\bm{\mathsf{x}}}^{\prime}(\omega^{\prime})\supseteq{\bm{\mathsf{x}}}(\omega^{\prime})$. Hence, ${\bm{\mathsf{x}}}^{\prime}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{x}}}$ by Axiom 4(a). Thus, by definition of $\geq_{\bm{\mathsf{X}}}$, $\omega\in D_{{\bm{\mathsf{x}}}^{\prime}}$ and $W_{\omega}\supseteq{\bm{\mathsf{x}}}^{\prime}(\omega)\supseteq{\bm{\mathsf{x}}}(\omega)=W_{\omega}$, whence ${\bm{\mathsf{x}}}^{\prime}(\omega)={\bm{\mathsf{x}}}(\omega)$ which in turn implies ${\bm{\mathsf{x}}}^{\prime}={\bm{\mathsf{x}}}$ by order consistency, Axiom 4(a). Thus, ${\bm{\mathsf{x}}}(\omega^{\prime})={\bm{\mathsf{x}}}^{\prime}(\omega^{\prime})=W_{\omega^{\prime}}$.
(Ad Part 2):  $D=\Omega$ iff for all $\omega\in\Omega$, $W_{\omega}\in X$. In view of Axiom 2, the set of roots of $(F,\supseteq)$ is given by precisely all $W_{\omega}$, $\omega\in\Omega$. The claimed equivalence follows.
(Ad Part 3):  Suppose that $(F,\pi,{\bm{\mathsf{X}}})$ is order consistent and maximal, and such that $X\neq\emptyset$. It suffices to show the claim ()($\ast$)( ∗ ) that there is exactly one ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ whose image contains a root of $(F,\supseteq)$. Indeed, then, by Axiom 33(b), $D\subseteq D_{\bm{\mathsf{x}}}$ and ${\bm{\mathsf{x}}}(\omega)=W_{\omega}={\bm{\mathsf{x}}}_{0}(\omega)$ for all $\omega\in D$, and by Part 1 of this lemma, proven just before, $D=D_{\bm{\mathsf{x}}}$, whence ${\bm{\mathsf{x}}}_{0}={\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$.
Regarding the proof of ()($\ast$)( ∗ ), note that by hypothesis there is $x\in X$. As $W_{\pi(x)}\supseteq x$, $W_{\pi(x)}\in X$. By Axiom 33(b), there is ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ with $x\in\operatorname{im}{\bm{\mathsf{x}}}$. Hence, if ()($\ast$)( ∗ ) did not hold true, there would have to be at least two distinct random moves ${\bm{\mathsf{x}}}_{1},{\bm{\mathsf{x}}}_{2}\in{\bm{\mathsf{X}}}$ whose images contain roots of $F$. By Part 1, the images of both ${\bm{\mathsf{x}}}_{1}$ and ${\bm{\mathsf{x}}}_{2}$ then would contain roots only. As ${\bm{\mathsf{x}}}_{1}\neq{\bm{\mathsf{x}}}_{2}$, by Axiom 4(a), we would have $D_{{\bm{\mathsf{x}}}_{1}}\cap D_{{\bm{\mathsf{x}}}_{2}}=\emptyset$. Then, let $\bar{\bm{\mathsf{x}}}={\bm{\mathsf{x}}}_{1}\cup{\bm{\mathsf{x}}}_{2}$, [^20] That is, $\bar{\bm{\mathsf{x}}}\colon D_{{\bm{\mathsf{x}}}_{1}}\cup D_{{\bm{\mathsf{x}}}_{2}}\to X$ such that $\bar{\bm{\mathsf{x}}}(\omega)={\bm{\mathsf{x}}}_{k}(\omega)$ for $\omega\in D_{{\bm{\mathsf{x}}}_{k}}$, for both $k=0,1$. and $\bar{\bm{\mathsf{X}}}=({\bm{\mathsf{X}}}\setminus\{{\bm{\mathsf{x}}}_{1},{\bm{\mathsf{x}}}_{2}\})\cup\{\bar{\bm{\mathsf{x}}}\}$.
Clearly, $D_{\bar{\bm{\mathsf{x}}}}=D_{{\bm{\mathsf{x}}}_{1}}\cup D_{{\bm{\mathsf{x}}}_{2}}$ would be a non-empty element of $\mathscr{E}$ and $\bar{\bm{\mathsf{x}}}$ would be a section of $\pi$ on $D_{\bar{\bm{\mathsf{x}}}}$. It is evident that $\bar{\bm{\mathsf{X}}}$ would be order consistent and induce a covering of $X$. By construction, ${\bm{\mathsf{X}}}$ would refine $\bar{\bm{\mathsf{X}}}$. Thus, maximality (see Axiom 4(c)) would imply the false statement $\bar{\bm{\mathsf{X}}}={\bm{\mathsf{X}}}$. Hence, ()($\ast$)( ∗ ) obtains and the proof is complete. ∎
@@
(Ad “order embedding”):  Let $({\bm{\mathsf{y}}}_{1},\omega_{1}),({\bm{\mathsf{y}}}_{2},\omega_{2})\in{\bm{\mathsf{T}}}\bullet\Omega$. If both ${\bm{\mathsf{y}}}_{1},{\bm{\mathsf{y}}}_{2}$ are elements of ${\bm{\mathsf{X}}}$, the implication $\Rightarrow$” is a consequence of the definition of $\geq_{\bm{\mathsf{X}}}$, and the other implication $\Leftarrow$” can be shown as follows: If ${\bm{\mathsf{y}}}_{1}(\omega_{1})\supseteq{\bm{\mathsf{y}}}_{2}(\omega_{2})$, then both moves belong to the same connected component of the rooted forest $(F,\supseteq)$ (whose existence Axiom 1 ensures). By Axioms 2 and 33(a), $\omega_{1}=\omega_{2}$. Then, Axiom 4(a) implies that ${\bm{\mathsf{y}}}_{1}\geq_{\bm{\mathsf{X}}}{\bm{\mathsf{y}}}_{2}$, whence ${\bm{\mathsf{y}}}_{1}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{2}$.
@@
If $k=2$, then ${\bm{\mathsf{y}}}_{1}(\omega_{1})\supseteq{\bm{\mathsf{y}}}_{2}(\omega_{2})$ is equivalent to the conjunction of these three statements: $\omega_{1}=\omega_{2}$, $\omega_{2}\in D_{{\bm{\mathsf{y}}}_{1}}$, and $w_{2}\in{\bm{\mathsf{y}}}_{1}(\omega_{1})$. By definition of $\geq_{\bm{\mathsf{T}}}$, this is equivalent to ${\bm{\mathsf{y}}}_{1}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{2}$ and $\omega_{1}=\omega_{2}$.
(Ad “bijection”):  Let $({\bm{\mathsf{y}}}_{1},\omega_{1}),({\bm{\mathsf{y}}}_{2},\omega_{2})\in{\bm{\mathsf{T}}}\bullet\Omega$ such that ${\bm{\mathsf{y}}}_{1}(\omega_{1})={\bm{\mathsf{y}}}_{2}(\omega_{2})$. As the evaluation map is an order embedding, which we have proven just before, we obtain ${\bm{\mathsf{y}}}_{1}={\bm{\mathsf{y}}}_{2}$ and $\omega_{1}=\omega_{2}$. Regarding surjectivity, let $y\in F$. If $y$ is terminal, then there is $w\in W$ with $y=\{w\}$ because $F$ is a decision forest on $W$ [^21] Although this is discussed in [3], we give an argument here for the reader’s convenience. Let $f$ be the map from Definition 1.3. If $y$ is terminal, then $(\mathcal{P}f)(y)=W(y)$ contains exactly one element, namely $\uparrow y$. As $f$ is injective, $y$ must be a singleton. Hence, $y={\bm{\mathsf{y}}}(\omega)$ for $\omega=\pi(y)$ and the random terminal node ${\bm{\mathsf{y}}}=\{(\omega,\{w\})\}$. If $y$ is a move, then, by the definition of ${\bm{\mathsf{X}}}$, Property 33(b), there is ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$ and $\omega\in D_{\bm{\mathsf{x}}}$ such that ${\bm{\mathsf{x}}}(\omega)=y$. ∎
@@
(Ad “forest”):  Let ${\bm{\mathsf{y}}}_{0}\in{\bm{\mathsf{T}}}$ and ${\bm{\mathsf{y}}}_{1},{\bm{\mathsf{y}}}_{2}\in\uparrow{\bm{\mathsf{y}}}_{0}$. As $D_{{\bm{\mathsf{y}}}_{0}}\neq\emptyset$ by Definition 2.9 and the definition of random terminal nodes, there is $\omega\in D_{{\bm{\mathsf{y}}}_{0}}$, and we have $\omega\in D_{{\bm{\mathsf{y}}}_{k}}$ and ${\bm{\mathsf{y}}}_{k}(\omega)\supseteq{\bm{\mathsf{y}}}_{0}(\omega)$ for both $k=1,2$. As $(F,\supseteq)$ is a forest, $\uparrow{\bm{\mathsf{y}}}_{0}(\omega)$ is a chain, thus ${\bm{\mathsf{y}}}_{1}(\omega)\supseteq{\bm{\mathsf{y}}}_{2}(\omega)$ or ${\bm{\mathsf{y}}}_{2}(\omega)\supseteq{\bm{\mathsf{y}}}_{1}(\omega)$. Hence, by Proposition 2.12, we get ${\bm{\mathsf{y}}}_{1}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{2}$ or ${\bm{\mathsf{y}}}_{2}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{1}$. We conclude that $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ is a forest.
@@
(Ad “decision forest”):  Let ${\bm{\mathsf{y}}}_{1},{\bm{\mathsf{y}}}_{2}\in{\bm{\mathsf{T}}}$ be such that ${\bm{\mathsf{y}}}_{1}\neq{\bm{\mathsf{y}}}_{2}$. If they are not comparable by $\geq_{\bm{\mathsf{T}}}$, that is neither ${\bm{\mathsf{y}}}_{1}>_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{2}$ nor ${\bm{\mathsf{y}}}_{2}>_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{1}$, then any maximal chain ${\bm{\mathsf{w}}}$ in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ with ${\bm{\mathsf{y}}}_{1}\in{\bm{\mathsf{w}}}$ satisfies ${\bm{\mathsf{y}}}_{2}\notin{\bm{\mathsf{w}}}$. Using the axiom of choice in the form of the Hausdorff maximality principle, there exists indeed such a ${\bm{\mathsf{w}}}$.
@@
For symmetry reasons, it remains to consider the case in that we have ${\bm{\mathsf{y}}}_{1}>_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{2}$. Then, by Proposition 2.12, we have for all $\omega^{\prime}\in D_{{\bm{\mathsf{y}}}_{2}}$, ${\bm{\mathsf{y}}}_{1}(\omega^{\prime})\supsetneq{\bm{\mathsf{y}}}_{2}(\omega^{\prime})$. As $D_{{\bm{\mathsf{y}}}_{2}}\neq\emptyset$, we can select $\omega\in D_{{\bm{\mathsf{y}}}_{2}}$. We can also select $w\in{\bm{\mathsf{y}}}_{1}(\omega)\setminus{\bm{\mathsf{y}}}_{2}(\omega)$. In particular, ${\bm{\mathsf{y}}}_{2}(\omega)\notin\uparrow\{w\}$. As the latter set is a maximal chain in $(F,\supseteq)$, there is $y_{3}\in\uparrow\{w\}$ such that $y_{3}$ and ${\bm{\mathsf{y}}}_{2}(\omega)$ cannot be compared by $\supseteq$. As ${\bm{\mathsf{y}}}_{1}(\omega)$ is also an element of the chain $\uparrow\{w\}$, this implies ${\bm{\mathsf{y}}}_{1}(\omega)\supsetneq y_{3}$. By Proposition 2.12, there is ${\bm{\mathsf{y}}}_{3}\in{\bm{\mathsf{T}}}$ such that $\omega\in D_{{\bm{\mathsf{y}}}_{3}}$ and $y_{3}={\bm{\mathsf{y}}}_{3}(\omega)$. This lemma also implies that ${\bm{\mathsf{y}}}_{2}$ and ${\bm{\mathsf{y}}}_{3}$ are not comparable via $\geq_{\bm{\mathsf{T}}}$. Hence, as we have proven just above, there is a maximal chain ${\bm{\mathsf{w}}}$ in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ with ${\bm{\mathsf{y}}}_{2}\notin{\bm{\mathsf{w}}}$ and ${\bm{\mathsf{y}}}_{3}\in{\bm{\mathsf{w}}}$. By Proposition 2.12, we obtain ${\bm{\mathsf{y}}}_{1}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}_{3}$, and using the fact that $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ is a forest, as already proven, we infer that ${\bm{\mathsf{y}}}_{1}\in{\bm{\mathsf{w}}}$. Hence there is a maximal chain in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ separating ${\bm{\mathsf{y}}}_{1}$ and ${\bm{\mathsf{y}}}_{2}$.
(Ad “rooted tree”):  In view of the preceding results, it suffices to show that $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ has a maximum. As $(F,\supseteq)$ is non-empty (as a rooted forest) and non-trivial (by sure non-triviality and non-emptiness), $X\neq\emptyset$. By Lemma 2.11, Parts 2 and 3, the map ${\bm{\mathsf{x}}}_{0}\colon\Omega\to X,\,\omega\to W_{\omega}$ is a random move, i.e. ${\bm{\mathsf{x}}}_{0}\in{\bm{\mathsf{X}}}$. Then, we have ${\bm{\mathsf{x}}}_{0}\geq_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}$ for all ${\bm{\mathsf{y}}}\in{\bm{\mathsf{T}}}$. Indeed, for all ${\bm{\mathsf{y}}}\in{\bm{\mathsf{T}}}$ and $\omega\in D_{\bm{\mathsf{y}}}$, $W_{\omega}$ is a root of the tree $(T_{\omega},\supseteq)$, whence $W_{\omega}\supseteq{\bm{\mathsf{y}}}(\omega)$.
@@
(Ad ${\bm{\mathsf{X}}}$ = set of moves):  Let ${\bm{\mathsf{x}}}\in{\bm{\mathsf{X}}}$. Then there is $\omega\in D_{\bm{\mathsf{x}}}$ and, by Axiom 33(a), ${\bm{\mathsf{x}}}(\omega)\in X$. Hence, there is $y\in F$ with ${\bm{\mathsf{x}}}(\omega)\supsetneq y$. By Proposition 2.12, there is unique ${\bm{\mathsf{y}}}\in{\bm{\mathsf{T}}}$ with $\omega\in D_{\bm{\mathsf{y}}}$ and $y={\bm{\mathsf{y}}}(\omega)$, and whence, by the same proposition, ${\bm{\mathsf{x}}}>_{\bm{\mathsf{T}}}{\bm{\mathsf{y}}}$. Hence, ${\bm{\mathsf{x}}}$ is a move in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$.
@@
On the other hand, the elements of ${\bm{\mathsf{T}}}\setminus{\bm{\mathsf{X}}}$, i.e. the random terminal nodes, are terminal nodes in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$ directly by definition of $\geq_{\bm{\mathsf{T}}}$. Hence, ${\bm{\mathsf{X}}}$ is the set of moves in $({\bm{\mathsf{T}}},\geq_{\bm{\mathsf{T}}})$. ∎
@@
$(\Omega,\mathscr{E})$ is a measurable space. $F$ is clearly a $W$-poset and the map assigning to any $w\in W$ the chain $\uparrow\{w\}$ is easily seen to be a bijection between $W$ and the set of maximal chains in $(F,\supseteq)$. Moreover, any chain in $(F,\supseteq)$ contains a maximal element, hence, by [3, Theorem 3], expressed in the language of the present paper, $F$ a decision forest on $W$ (Axiom 1).
@@
For all $k=0,1,2$, ${\bm{\mathsf{x}}}_{k}$ is defined on $\Omega$ which is an event. Clearly, ${\bm{\mathsf{x}}}_{k}$ is a section of $\pi$. The union of the images of ${\bm{\mathsf{X}}}$’s elements is the set of nodes with at least two elements. As $\{\{w\}\mid w\in W\}\subseteq F$, this union equals $X$. Thus Axioms 33(a) and 33(b) are satisfied.